# Robust surface superconductivity and vortex lattice in the Weyl semimetal -PtBi2

### v1

@@ Acknowledgments
We acknowledge support by the Spanish Research State Agency (PID2023-150148OB-I00, TED2021-130546BI00, PDC2021-121086-I00 and CEX2023-001316-M), the European Research Council PNICTEYES through grant agreement 679080 and VectorFieldImaging Grant Agreement 101069239, the EU through grant agreement No. 871106 and by the Comunidad de Madrid through projects projects TEC-2024/TEC-380 “Mag4TIC”. We acknowledge the “QUASURF” project [SI4/PJI/2024-00199] funded by the Communidad de Madrid through the direct grant agreement for the promotion and development of research and technology transfer at the Universidad Autónoma de Madrid. We have benefited from collaborations through EU program Cost CA21144 (superqumap), and from SEGAINVEX at UAM in the design and construction of STM and cryogenic equipment. Work done at Ames National Laboratory (PCC, SLB, LLW, ZL) was supported by the U.S. Department of Energy, Office of Basic Energy Science, Division of Materials Sciences and Engineering. Ames National Laboratory is operated for the U.S. Department of Energy by Iowa State University under Contract No. DE-AC02-07CH11358.
Supplementary Material of: Robust surface superconductivity and vortex lattice in the Weyl semimetal -PtBi2
I Single crystal growth of PtBi2
@@
II Comparison with the superconducting properties found in literature
Bulk resistivity measurements have found a transition to the superconducting state at 0.6 K-1.1 K[18, 13] that can be enhanced up to K by doping with Rh[18] or by hydrostatic pressure[23]. Thin flakes with a thickness of the order of tens of nanometers have shown a Berezinskii-Kosterlitz-Thouless transition at mK,[22]. Techniques probing the surface have found a plethora of different critical temperatures ranging from K (point-contact) [14], to K (SQUID on tip) [15] and to 8-14 K [12] (ARPES). Most remarkably, STM measurements found superconducting gaps compatible with a of the order of 80 K [15] and a T. However, superconducting properties probed in Ref. [15] were extremely inhomogeneous and even contained non-superconducting regions [17]. Furthermore, the vortex lattice could not be observed in Ref. [15] despite having performed measurements under high magnetic fields. Other work in -PtBi2 did not find evidence for superconductivity in STM [16] and ARPES (down to 3 K) [20].
III Further aspects of the STM experiment
We cleaved samples in-situ at 4.2 K or below. Samples were cleaved by gluing a post of alumina on top of the samples and hitting the alumina by moving the sample holder below a beam. We measured hundreds of fields of view in three different samples, obtaining the behavior discussed in the main text in each of them.
We used tips of Pt-Ir, unless in the experiment with superconducting tip, where we used Pb tips. To prepare and clean the tips we used the tip preparation methods described in Ref. [64, 65], which essentially consists of moving the tip towards a sample of the same material as the tip and making controlled repeated indentations.
We would like to remark that we usually find a single surface termination when measuring a certain sample of -PtBi2. To obtain data from the two surface terminations, we first cleaved one sample and measured it, finding atomically flat terraces with termination A (Fig. 1 of the main text). The cleaved piece was attached to a rope and was let to fall down to the bottom of the inner vacuum chamber of the dilution refrigerator. We then took this part and mounted it in such a way as to make sure that the upper surface was at the other side as compared with previous experiment. We then cooled again and cleaved the sample again. We found, in that new cleave, always atomically flat terraces with the same B-surface termination, opposite to the previous one (termination B in Fig. 1 of the main text). The shape and superconducting behavior on nanometer-sized flakes was the same for both surface terminations in atomically flat areas.
It is not fully clear why we do not observe atomic resolution on flakes, but instead a surface with a certain corrugation larger than atomic sizes. We speculate that the surface of the flakes is not completely flat. Interestingly, the flat areas on the flakes are sufficiently large so that surface states and superconducting vortices form.
During scanning, the current was most often around 1 nA and the bias voltage of a few mV. We usually present the tunneling conductance normalized to its value well above the superconducting gap edge. Tunneling conductance maps were obtained by cutting the feedback loop at each pixel on the scan area and making a full tunneling conductance vs bias voltage measurement. We also provide DFT calculations, performed as described with detail in the in the next section.
IV DFT calculations

### v2

@@ Acknowledgments
We acknowledge discussions with A. Kaminski, Miguel Águeda and with Isidoro Poveda. We acknowledge support by the Spanish Research State Agency (PID2020-114071RB-I00, PID2023-150148OB-I00, TED2021-130546BI00, PDC2021-121086-I00 and CEX2023-001316-M), the European Research Council PNICTEYES through grant agreement 679080 and VectorFieldImaging Grant Agreement 101069239, the EU through grant agreement No. 871106 and by the Comunidad de Madrid through projects TEC-2024/TEC-380 “Mag4TIC”. We acknowledge the “QUASURF” project [SI4/PJI/2024-00199] funded by the Comunidad de Madrid through the direct grant agreement for the promotion and development of research and technology transfer at the Universidad Autónoma de Madrid. We have benefited from collaborations through EU program Cost CA21144 (superqumap), and from SEGAINVEX at UAM in the design and construction of STM and cryogenic equipment. Work done at Ames National Laboratory (PCC, SLB, LLW, ZL) was supported by the U.S. Department of Energy, Office of Basic Energy Science, Division of Materials Sciences and Engineering. Ames National Laboratory is operated for the U.S. Department of Energy by Iowa State University under Contract No. DE-AC02-07CH11358. First-principles calculations have used resources of the National Energy Research Scientific Computing Center (NERSC), a DOE Office of Science User Facility.
I Supplemental Material
II Superconducting properties in different fields of view
II.1 Characterization of the corrugated surface areas on flakes
As we show in the main text, we find areas which are atomically flat, and also areas with a certain corrugation. The corrugated areas are located on top of the atomically flat areas, and we name these in the following nanometer-sized flakes.
We have imaged the surface of our cleaved samples after STM measurements using high-resolution SEM to characterize the origin and composition of the nanometer-sized flakes. We have used an Electron Beam Lithography eLINE-PLUS from Raith GmbH, which incorporates a Bruker XFlash 6I30 EDS detector. In Fig. S1 we show several SEM images of different sizes. We can see on the edge of the sample (Fig. S1(a)) that the step edges often have 120∘ corners, reflecting the trigonal symmetry of the crystal lattice. Furthermore we find flakes close to the edges. These flakes have probably been caused during cleaving and we can find them all over the surface, including flat areas well inside the sample (Fig. S1 (b)). The flakes have sizes similar to those that we have observed in our STM measurements. In Fig. S1(c) we show a high resolution SEM image on the area marked by an ocre rectangle in Fig. S1(b). We see that all flakes have borders with an hexagonal shape, marked in Fig. S1(c) by purple dashed lines. However, their surface is not completely flat, contrasting the surface of the rest of the sample. There are flakes with a large surface corrugation (upper right part of Fig. S1(c)) and others with much less corrugation (left part of Fig. S1(c)). This agrees remarkably well with our STM observations—flakes that are one or a few unit cells high and are relatively flat (as those discussed below in Fig. S6(a,c,d) and in Fig. S9(d,g)), and flakes showing surfaces with enhanced corrugation.
To analyze the elemental composition, we have performed Energy Dispersive electron Spectroscopy (EDS) measurements on some of these flakes (sites 1-3 marked in in Fig. S1(b)) and on top of the terrace (site 4 in Fig. S1(b)). Results are shown in Fig. S1 (d). There is no difference between spectra performed on the flakes (sites 1-3) and on flat terraces (site 4). A standardless quantitative analysis yields ratios of Pt/Bi atomic concentration of 0.46 (site 1), 0.49 (site 2), 0.46 (site 3) and 0.46 (site 4). Within the experimental uncertainty of 0.05 in the Pt/Bi ratio, this coincides with the stoichiometric ratio of PtBi2. We note that the electron–sample interaction volume may introduce a small uncertainty in these measurements. From the measurement parameters (working distance = 10.3 mm, accelerating voltage = 15 kV), we estimate an electron beam radius of approximately 400 nm and an interaction depth of about 700 nm. The lateral beam size is sufficiently small that the beam falls entirely within most flakes. However, the interaction depth implies that, for sites 1–3, the EDS signal is expected to include a contribution from the underlying flat terrace beneath the flakes. Nevertheless, if these flakes were a species different from PtBi2, we would expect the corresponding EDS spectra to differ at least qualitatively from that measured at site 4.
II.2 Superconducting properties in nanometer-sized flakes, compared to those in flat areas
We show topographies and conductance maps in fields of view that contain atomically flat surfaces and nanometer-sized flakes in Fig. S2, as a function of magnetic field, and in Fig. S3 in a different field of view. Topographies are shown in Fig. S2(a) and Fig. S3(c), and the corresponding zero-bias conductance map are shown in Figs. S2(b-d) and Fig. S3(a). In Figs. S2(f-h), and Figs. S3(b,e) we show the conductance profiles crossing the boundary between flat surfaces and nanometer-sized flakes. Nanometer-sized flakes present a superconducting density of states which is equivalent to the one observed in atomically flat surfaces. We observe the same gap size, and coherence peaks with the same shape at zero field and under magnetic fields.
III Vortex lattice observations with STM
III.1 Tip-vortex interaction
We illustrate schematically the possible interaction between tip and vortex lattice in Fig. S4. As is well known, a STM tip is most often blunt, but there is always a small atomic size apex which dominates in the tunneling process, because of the exponential dependence of the tunneling current with distance [71]. We represent the tunneling apex by a small triangular-like tip in Fig. S4. The tunneling current from this apex is shown schematically in red. The overall shape of the tip is blunt and we represent it by a semi-spherical object with a radius of several nm. The electrostatic interaction between tip and sample occurs over a much larger scale than tunneling. This length scale depends on the size of the tip but is at least several nm in diameter [72, 73, 74, 75, 76, 77]. The electrostatic interaction between tip and dipoles on the sample contains several components, such as image charges and the actual dipolar interaction [72, 75, 76]. The latter can be, in principle, attractive and repulsive, depending for instance on the sign of the bias. Independently on the sign of the interaction, for a certain electrostatic configuration, there is an equilibrium distribution of vortices below the tip, most often consisting of the vortex lattice below the capacitor built by the spherical tip (with nm or tens of nm radius). The charge of these vortices can eventually influence the electrostatic tip-sample interaction, particularly if for two-dimensional vortices.
As shown in Ref. [53], the differences in the chemical potential between superconducting and normal quasiparticles induce a charge redistribution in vortices. Metallic screening does not fully eliminate the charge. This results in an electric dipole oriented perpendicular to the surface whose strength is of the order of where is the elementary charge and Å(which gives D, one Debye). Such a minute electric dipole has been often considered previously and has an influence in several properties of superconductors, such as the Hall effect in the mixed state [78, 79, 80, 81, 82]. To estimate the force generated by the electric field of the tip on the vortex, we first consider the corresponding interaction energy, . The dipole sits on the vortex core, but its field extends over a distance of order of the superconducting coherence length . We consider a tip which is at a distance smaller than and is blunt and large in size. The effective electric field is then , where is the voltage applied to the tip. We can then estimate the force acting on the vortex as . For V, nm we obtain fN.
The pinning force can be estimated by considering the free energy per unit volume from the thermodynamic critical field and the vortex volume [83]. Due to the tubular structure of vortices, the pinning force depends critically on the vortex length. Estimates of the pinning force per unit vortex length are often of order of N/m [83, 84]. In PtBi2, the bulk is normal, which implies that superconductivity is restricted to a small layer close to the surface and vortices are essentially two-dimensional pancakes. Taking a conservative estimate for the thickness for the superconducting layer of 1 nm (approximately two unit cells along the c-axis, unit cell size along the c-axis is of 0.617 nm), we estimate fN [85], which is one order of magnitude below our estimate of .
To delve into this aspect, we can estimate the possible shift of vortex positions by considering collective pinning in two dimensions [86]. There is a region where the vortex lattice is correlated, of size . The elastic energy density is then given by in three dimensions, where is the elastic shear modulus. In two dimensions, the energy is independent on , and is given by , where is the thickness of the superconducting layer. We can estimate using [87], with the thermodynamic critical field and . To relate the force with , we can use . For magnetic fields at about half the upper critical field and for forces in the fN range we find a vortex shift which is of a few times the coherence length . Therefore, the two-dimensional vortex lattice is very soft.
On flakes, the vortex pinning increases. We have observed flakes above 5 nm in height, roughly an order of magnitude larger than the c-axis unit cell. The eventual increase in vortex length leads to an increased pinning force up to . For example, in flakes that are close to a c-axis unit cell size, vortices are still unpinned but remain static for flakes with larger thickness, as we show below. This is schematically represented in Fig. S4(d) where we show an increase in the effective length of the magnetic vortex. This increases the pinning force, eventually leading to . In that case, we have the situation shown in Fig. S4(e,f). The vortex lattice on the flakes remains fixed (dark blue on Fig. S4(e,f), ), but vortices on the flat two-dimensional surface are moved by the interaction with the tip (light blue on Fig. S4(e,f), ).
In practice, the interaction can just lead to changes in the position of a few vortices, because the hexagonal vortex lattice is not perfect and the vortex positions are slightly influenced by small defects. With very weak pinning, we obtain the result shown schematically in Fig. S4(b,c). When we move the tip the equilibrium position of some vortices move together with the tip in such a way as to maintain the distribution of charges below the tip.
Thus, we conclude that on large and flat surfaces, , vortices are highly mobile and remain undetected. However, on flakes of sufficient thickness, , the vortex lattice is pinned and STM reveals the presence of vortices. For comparison, the pinning strength, estimated by taking , with the thickness of the superconducting layer, is of K for nm, similar to the weak pinning observed in trilayer graphene and below the one observed in disordered thin films [24, 27]. The enhanced mobility of vortices on atomically flat surfaces is further evidenced in Fig. 3(g,h) of the main text where we find that conductances curves on atomically flat surfaces are equivalent to those taken on flakes. We provide experimental evidence for the thickness-dependent pinning of vortices and tip-induced vortex motion on nanometer-sized flakes further below.
Absence of vortices has been reported previously in highly disordered and very thin films particularly close to the superconductor to insulator transition [88, 89, 90]. An enhanced vortex mobility and tip-sample interactions could play a certain role in those observations too, as well as in other two-dimensional superconducting systems [22, 23, 24, 25, 26, 27, 28, 29, 30].
III.2 Temperature dependence of the vortex lattice
We show the effect of increasing temperature on the vortex lattice of PtBi2 in Fig. S5(a-d). The scans are made consecutively and in the same field of view. The nanometer-sized flake slightly changes in lateral size with subsequent scans, possibly due to tip-sample interactions (see Fig. S9). However, the surface remains large enough to identify the vortex lattice at the applied magnetic field. The hexagonal vortex lattice remains distinguishable up to 1.2 K, but is more disordered as compared to the vortex lattice at lower temperatures. At 1.5 K, there are practically no traces of superconducting vortices. However, the average zero bias tunneling conductance normalized to its value at high voltages within the flake is well below one. This suggests that the flake is still superconducting, but with vortices strongly fluctuating their position, giving a time averaged tunneling conductance which is neither one, as observed in a pinned vortex core, nor very low, as observed in between vortices [54, 88, 90]. At 2 K, the zero bias tunneling conductance on the nanometer-sized flake is homogeneous, suggesting strongly thermally activated vortex motion, and a melted vortex lattice [54, 56, 55, 90, 27]. Remarkably, the tunneling conductance in the flat areas outside the nanometer-sized flake strongly increases with temperature too. This suggests that the thermal jitter of vortices induces a finite tunneling conductance also on the flat areas.
III.3 Vortex behaviour with magnetic field history
The vortex lattice on flakes is observed with similar properties when reversing the direction of the magnetic field. Furthermore, zero field cooled and field cooled experiments lead to similar observations too. This is to be expected due to the strong field penetration in two-dimensional superconductors.
III.4 Thickness dependence
Of interest is also a closer analysis of the field of view shown in Fig. 3(c,d) of the main text. In Fig. S6(a) we show the topography acquired simultaneously with the tunneling conductance map shown in Fig. 3(c) of the main text. In the center of the field of view there is a large nanometer-sized flake. On the bottom left corner (orange square) there is a stripe-like flake separating two atomically flat regions. In Fig. S6(b) we show the tunneling conductance map corresponding to this area. We see that individual vortices are barely observed, and instead there is a central line with a large density of states. It is important to note that this striple-like nanometer-sized flake is smaller in height than the one present on the center of Fig. S6(a). The height profiles, shown in Fig. S6(c,d) show a size difference of approximately an order of magnitude. The height of the one-dimensional part of the flake is roughly one unit cell along the c-axis (0.619 nm). This implies that the vortex length is smaller on the stripe-like nanometer-sized flake, which allows for vortex mobility and produces a one-dimensional stripe of mobile vortices.
Thus, there is a gradual evolution with increasing flake height from the highly mobile vortices on flat surfaces to the completely pinned vortices on large nanometer-sized flakes. This suggests that vortices on flakes are longer than those on the atomically flat areas, enhancing their pinning, similar to pancake vortices in cuprate superconductors [91, 92]. We also note that vortices seem to generally align with the edges of the flakes. The vortex Fig. S3(a) is located at an edge of the nanometer-sized flake. Furthermore, the vortex lattices shown in Fig. 3(b,c,d) of the main text and in Fig. S5 are oriented following the edges of the flakes.
III.5 Tip-vortex interaction on nanometer-sized flakes.
In Fig. S7 we show results on a field of view containing a large single flake with an applied magnetic field of 0.2 T, which is relatively low (far from ). This image has been taken with a tunneling current of about 2 nA. Both features, increased tunneling current (decreased tip-sample distance) and low field, provide a stronger tip-sample interaction. In Fig. S7(a,b) we show the topography and a line scan. In Fig. S7(c,d) we show two zero bias conductance images taken in the same field of view, (d) taken immediately after (c). White dots in Fig. S7(c,d) provide the vortex positions in each scan. In Fig. S7(d) we show as grey points vortex positions of Fig. S7(c). We see that vortices are slightly displaced, indicating a strong tip-sample interaction. This is further evidenced in Fig. S8 where we present a situation with several nanometer-sized flakes at 0.15 T. In this situation we have studied the vortex configuration as a function of decreasing tip-sample distance. We have increased the tunneling conductance from 0.008 to 0.40, i.e. by two orders of magnitude ( is the quantum of conductance). This corresponds to a decrease in tip-sample distance of about Å. We stop just before the contact between tip and sample (expected for a tunneling conductance [93]). In Fig. S8 (a-c) we show the topography and two line scans across flakes and atomically flat surfaces. In Fig. S8(d-g) we show zero bias conductance images taken in the same field of view with increasing tunneling conductance. For low tunneling conductance, there is only one vortex per flake as seen in S8 (d). When we increase tunneling conductance we find in Fig. S8 (e) that vortices are particularly unstable and are easily moved by the tip: vortices appear “cut” horizontally along the direction of the scan. When further increasing the tunneling conductance, additional vortices appear on the flakes (see vortices marked with red arrows in Fig. S8(f,g)). In Fig. S8 (g), when tip-sample distance is minimal, there are roughly two vortices per flake arranged in two “rows”. The second vortex row has been dragged by the tip from atomically flat surfaces onto the nanometer-sized flakes due to the increased tip-vortex interaction.
The evolution of the vortex configuration shown in Figs. S7,S8, together with the thickness-dependent pinning in Fig. S6, shows that the pinning force acting on nanometer-sized flakes is of the same order as the electrostatic interaction between STM tip and vortices, . By reducing the distance between tip and sample at working at a low applied magnetic field, we have been able to partially overcome the pinning and drag vortices into a new configuration.
III.6 Interaction between the vortex lattice and nanometer-sized flakes
Another example of an atomically flat surface at an applied magnetic field of 0.5 T is shown in Fig. S9(a). This surface presents a “hole” on the upper left corner, with an approximate depth of 10 nm. In Fig. S9(b) we see the result of a first scan. Inside the hole we find a nanometer-sized flake. Although the superconducting gap is perfectly homogeneous over the whole surface, we see the vortices on the nanometer-sized flake within the hole Fig. S9(c). After making several scans, we find the topography shown in Fig. S9(e). The surface inside the hole is now perfectly flat, because interactions with the tip have modified the position of the nanometer-sized flake, possibly removing it altogether from this field of view. We then see no trace of the vortices in Fig. S9(f) and instead a spatially homogeneous superconducting tunneling conductance with the properties discussed in the main text.
IV Upper critical field and vortex core size as a function of the magnetic field
In Fig. S10 we show the normalized tunneling conductance curves as a function of the magnetic field, taken in between vortices on top of flakes. We find suppressed coherence peaks and a gradual filling of the gap when increasing the magnetic field. To estimate the value of the upper critical field we trace the zero bias conductance as a function of the magnetic field (inset of Fig. S10) and find that it reaches the normal state value at a magnetic field of approximately T.
The magnetic profile of vortices expands in a two-dimensional superconductor [59]. But the spatial profile of the superconducting order parameter around the vortex core does not depend on the dimensionality and remains of the order of the coherence length perpendicular to the magnetic field, . There is a characteristic evolution of the profile of the superconducting density of states around a vortex with magnetic field. When the vortex density increases with magnetic field, the slope of the superconducting order parameter , given by eventually presents a magnetic field dependence which was considered in Ref. [95, 94]. , with being the so-called vortex core size which coincides exactly at the upper critical field with . It was shown that the radial dependence of the tunneling conductance can be written as:
where , , and is half the intervortex distance. The model relates the density of states with the tunneling conductance using deGennes model [96] and assumes a smooth overlap on the order parameter value between different vortices at . In Fig. S11 we show the vortex core size , obtained from fitting the radially-averaged tunneling conductance along vortex cores to equ. 1. We find that the vortex core size is slightly decreasing with the magnetic field and tends towards the coherence length nm at the critical field T. This behavior is expected for a superconductor with an electronic mean free path of order of [94].
V Josephson coupling between -PtBi2 and superconducting tips
To obtain further evidence of the robustness of the superconducting properties of -PtBi2 we performed controlled indentations of the tip into the sample, picking up pieces of PtBi2. We observed the Josephson effect and traced its evolution when changing the tunneling current. We observed similar behavior as found previously in systems such as Pb or Al [97]. There are peaks in the tunneling conductance at twice the superconducting gap and at the superconducting gap, due to Andreev reflection. In addition, there is a pronounced peak at zero bias which evidences the passage of a Cooper pair (Josephson) supercurrent (Fig. S12 (a)). When measuring with a tip of Pb, we also observe a Josephson peak (Fig. S12 (b)) and the features in the tunneling conductance characteristic of tunneling between superconductors with different gap sizes [98], marked by arrows in Fig. S12 (b).
VI Further aspects of the STM experiment
We cleaved samples in-situ at 4.2 K or below. Samples were cleaved by gluing a post of alumina on top of the samples and hitting the alumina by moving the sample holder below a beam. We measured hundreds of fields of view in four different samples, obtaining the behavior discussed in the main text in each of them.
We used tips of Pt-Ir, except in the experiment with superconducting tip, where we used Pb tips. To prepare and clean the tips we used the tip preparation methods described in Ref. [97, 93].
We would like to remark that we usually find a single surface termination when measuring a certain sample of -PtBi2. To obtain data from the two surface terminations, we first cleaved one sample and measured it, finding atomically flat terraces with termination A (Fig. 1 (b) of the main text). The cleaved piece was attached to a rope and was let to fall down to the bottom of the inner vacuum chamber of the dilution refrigerator. We then recovered this part and mounted it in such a way as to make sure that the upper surface was at the other side as compared with previous experiment. We then cooled and cleaved the sample again. We found, in that new cleave, always atomically flat terraces presenting topographies as shown in Fig. 1(c) of the main text, as opposed to the topographies found in the original cleave (Fig. 1(b) of the main text). Therefore, these two types of images present the two types of surfaces obtained from a cleave, A and B in Fig. 1(b,c) of the main text. The superconducting behavior was the same for both surface terminations.
Flakes do not present large atomically flat surfaces, but instead inclined planes and height changes that are of atomic size or below, but do not form a periodic spatial pattern. This suggests that flakes are not completely flat and consist of many terraces (see also Scanning Electron Microscopy (SEM) results below). We note, however, that we can find flakes of one unit cell height, as shown in Fig. S6(a,d), and others with about eight unit cells, as shown in Fig. S9(d).
During scanning, the current was most often below 1 nA and the bias voltage of a few mV. We increased the tunneling current until 50 nA, observing the same features (absence of vortex lattice in atomically flat areas and presence in corrugated areas). We usually present the tunneling conductance normalized to its value well above the superconducting gap edge. Tunneling conductance maps were obtained by cutting the feedback loop at each pixel on the scan area and making a full tunneling conductance vs bias voltage measurement.
Sometimes it is useful to fit the tunneling conductance to a lifetime broadened density of states [99, 100]. This often allows discussing broadened quasiparticle peaks. In our case, this does not lead to appropriate fits, because the tunneling conductance deviates from zero very rapidly when increasing the finite lifetime in the model, whereas it remains very small in the experiment. Therefore, we cannot fit at the same time broadened quasiparticle peaks and the low energy tunneling conductance.
VII Single crystal growth of PtBi2 and bulk characterization
@@
We show measurements of the zero field cooled magnetization and the AC susceptibility as a function of temperature in our single crystals in Fig. S13. To maximize any possible trace of superconductivity in the magnetization, we used samples with a plate-like form, oriented perpendicular to the magnetic field (c-axis parallel to the magnetic field). As we can see in Fig. S13, there are no traces of superconductivity.
VIII DFT calculations